OpenStudy (anonymous):

Can anybody tell me 2y-3=13 please? :)

OpenStudy (anonymous):

2y-3 = 13 2y = 13+3 2y = 16 y = 16/2 y = 8

OpenStudy (anonymous):

Thankyou geniusy_2006 :) xxxxx